OpenStudy (anonymous):

can someone please factor the expression 48x^2+72x+15 or is it a prime polynomial

OpenStudy (anonymous):

First factor out a 3.

OpenStudy (anonymous):

\[48x ^{2}+72x+15\rightarrow 3(16x ^{2}+24x+5)\]

OpenStudy (anonymous):

Let me say it is not prime, it is factor-able, and you will need to use guess and check to figure out how to factor it.

OpenStudy (anonymous):

thank you for the tip i got it figured out

OpenStudy (anonymous):

heres another one i need help with \[5\sqrt{3}-\sqrt{98}+\sqrt{48}+6\sqrt{2}\]